Suppose you were to take a sample of water from a nearby lake. Determine the concentration of O2 in this water sample at 25 °C given that at sea level the Earth's atmosphere contains 77.79% N2 and 21.21% O2.
Answer #1

Partila pressure of O2 = atmospheric pressure * composition of O2 / 100 = 1 * 21.21 / 100 = 0.2121 atm

Using henry's law:

partial pressure / Concentration = KH

concentration = 0.2121 / 769.23 = 2.76 * 10-4 M